OpenStudy (anonymous):

Find the equation in standard form which passes through (3, –6) and (–2, –1). You must show all work.

OpenStudy (apoorvk):

Of a straight line ofcourse?

OpenStudy (anonymous):

yeah

OpenStudy (apoorvk):

\[\text{The equation of a straight line passing through two points A(}x_1,y_1\text{) and B(}x_2,y_2\text{) is:}\] \[\frac{y - y_1}{x - x_1}=\frac{y_2 - y_1}{x_2 - x_1}\] Whaddya think now? Can you do this?

OpenStudy (anonymous):

not a chance! hahahaha

OpenStudy (lgbasallote):

correction @apoorvk it's actually \[(x_2 - x_1)(y - y_1) = (y_2 - y_1)(x - x_1)\]

OpenStudy (apoorvk):

@lgbasallote - open your eyes and see what happens if I rearrange it :D

OpenStudy (lgbasallote):

OMMIGOSH! :O possibilities opened *_*

OpenStudy (apoorvk):

"Open thy eyes and open a world of new possibilities." - Some geek.